VT008A

PI: Gurvits
Title: VSOP test

Test: Jan 1997
Frequency: 6 cm
Mode: VLBA/VSOP
Obs time (UT): 13 EU 10:00(13/01)-14:40(13/01)
Stations: JbWbOnMcShTrScLaFd
Processor: Socc
EU Mk3 Tapes/St.: Europe 2 (thick), VLBA 1 (thin)

Comment


To give feedback on this experiment, click here

Experiment Feedback

Westerbork (Tony Foley at Mon Jan 13 07:45:36 EST 1997) Minor failures
Used telescope RT7 at position: 3828647.18 -443450.02 5064923.08 (this is ~144m east of RT6, used in December) System temp equivalent ~ 2400Jy Linear polarization NO phasecal (unit is broken)

Jodrell (Alastair Gunn at Tue Jan 14 05:35:06 EST 1997) Minor failures
Minor problem with bad scheduling. Did not reach source 1739+522 until 13:17:45 UT, thus missing scan 7 (tape 1). Did not reach source 1803+784 until 13:45:30 UT thus missing beginning of scan 9 (tape 2). Lost telescope for approximately one minute at about 13:30 UT. No other known problems. Weather overcast, dry, with occasional sunny periods, wind calm. Tapes JPL00872,5585 and RRL00078,C2F5 sent to Socorro via Fedex on Tuesday 14th 1997.

Medicina (Tiziana Venturi at Tue Jan 14 10:30:30 EST 1997) Minor failures
VC05 did not work for the first two hours of the experiment. The problem was spotted and fixed, and the experiment ran smoothly afterwards.

Onsala (John Conway at Tue Jan 14 11:24:41 EST 1997) Success
No Tsys from VC2 because after being replaced in ran had a negative tpzero offset. Otherwise no known problems